OpenStudy (anonymous):

solve the equation for x 2ln x= ln(2x-3) +ln(x-2) i know that the answer is 6 but i have no idea how to get it... can anyone help????

OpenStudy (freckles):

do you know the product rule? ln(a)+ln(b)=ln(ab) Use that rule on right hand side. for the left hand side, use the power rule. That is, r ln(x)=ln(x^r) Then once you have ln(u)=ln(v) set u=v and solve for x.

OpenStudy (freckles):

@kathyneedshelpplease are you there?

OpenStudy (anonymous):

thank you

OpenStudy (freckles):

well if you want you can post here your work and i will check it or if you need any other help with this problem please let me know